olve the problem.
In a right triangle, one acute angle is 54degree more than twice the other. Find each acute angle.

Answer by checkley71(8403)

90+(54+2X)+X=180 OR 90+54+3X=180 OR 144+3X=180 OR 3X=180-144 OR 3X=36 OR X=12 & 54+2*12=78 THIS 90+78+12=180 OR 180=180